OpenStudy (anonymous):

PLEASE i will give medal and fan :) The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent?

OpenStudy (anonymous):

The area of the base of the tank The area of the curved sides of the tank The volume of 3 of the same cone-shaped tanks The volume of 6 of the same cone-shaped tanks

OpenStudy (anonymous):

So which part of the question do you not understand? Do you understand how they got \[\frac{ 1 }{ 3 }\pi(x-7)^{2}(x-3)\]

OpenStudy (anonymous):

Basically, the factors represent the volume formula of a cone which is\[(\frac{ 1 }{ 3 })\pi*r ^{2}*h\]

OpenStudy (anonymous):

yeah i dont even know how they got that ..

OpenStudy (anonymous):

wait so the height 3?

jigglypuff314 (jigglypuff314):

hwtrain has it right :) they gave a volume and asked what it would represent without the (1 over 3) part and so the difference between (1/3)volume and volume is a multiple of 3

OpenStudy (anonymous):

so do i multiply 1/3 * 3 ?

jigglypuff314 (jigglypuff314):

more like which option has "3" in it? :P

OpenStudy (anonymous):

Lol i thought about that ... thank you !